OpenStudy (mellamokatie):

Write an equation in point-slope form for the line through the given point with the given slope. (8, 3); m = 6 A. y + 3 = 6(x-8) B. y - 3 = 6 (x -8) C. y - 3 = 6(x + 8) D. y + 3 = 6x + 8 @rebeccaxhawaii

Vocaloid (vocaloid):

use the template y - y0 = m(x-x0) where m = slope, x0 = your x-coordinate and y0 = your y-coordinate

OpenStudy (mellamokatie):

Hey I got to go I will be back tonight thanks for all the help today.

OpenStudy (mellamokatie):

What do you mean?

OpenStudy (mellamokatie):

@Vocaloid

OpenStudy (mjdennis):

How much do you know about the equation of a line.

OpenStudy (mellamokatie):

Nothing xD

OpenStudy (mjdennis):

I doubt it. You don't get to this kind of problem knowing nothing. What does slope _mean_? What's the slope in this problem?

OpenStudy (mellamokatie):

What? I dont know it...

OpenStudy (mjdennis):

Almost always, slope will be represented by 'm'. Two or three problems ago, @rebeccaxhawaii gave you the formula for the slope of a line. You even said you were writing it down.: http://openstudy.com/users/mellamokatie#/updates/57478ddfe4b064779b120444 You should also have learned that slope is a measure of how "steep" a line is.

OpenStudy (mellamokatie):

Yeah I wrote it down I have it here I didnt know how I would use that in the formula

OpenStudy (throneofbooks):

point slope form= y-y1=m(x-x1)

OpenStudy (mjdennis):

@Throneofbooks , were working on getting there.

OpenStudy (mellamokatie):

yeah

OpenStudy (mjdennis):

@MellamoKatie , sorry, I was trying to draw you a picture about slope. We can work with @Throneofbooks formula, though. One sec

OpenStudy (mellamokatie):

Oh okay thanks anyway

OpenStudy (mellamokatie):

Yes I know that so what would I plug in?

OpenStudy (mjdennis):

If we have two points (x1,y1) and (x2,y2), we just plug them in and find m, the slope. This is a basic formula, math people work out different equations from this one.

OpenStudy (mjdennis):

We have to change it up for your problem: One sec.

OpenStudy (mellamokatie):

oh okay

OpenStudy (mjdennis):

In your problem, we _knjow_ the slope m and 1 point. So I am going to change the LABELS in that equation, then show you where to put your numbers.

OpenStudy (mellamokatie):

Oh great that's what I need help with ! :D

OpenStudy (mjdennis):

The problem asks for "point-slope form". All we do is use algebra to move around the pieces of the formula. Here it is with variables instead of numbers:\[m = \frac{ (x - x _{1}) }{ (y - y _{1}) }\] \[m * (y - y _{1}) = (x - x _{1})\]

OpenStudy (mjdennis):

Shoot, I wrote your formula wrong. Go take a break, and I'll fix it

OpenStudy (mjdennis):

Slope is the change in y divided by the change in x. Like a protractor or a compass, it is a different way to measure the direction or angle a line makes. \[m = \frac{ (y _{2} - y _{1}) }{ (x _{2} - x _{1}) }\]

OpenStudy (mjdennis):

\[m = \frac{ (y _{2} - y _{1}) }{ (x _{2} - x _{1}) }\] \[m = \frac{ (y - y _{1}) }{ (x - x _{1}) }\] Now, x-without-subscript is a variable -- we don't know what it is, so we just leave it as 'x' and the same for 'y'. But we know 'm', and (x1,y1), right. Go ahead and put them in, and tell me your equation.

OpenStudy (mellamokatie):

Wait there is only (8, 3) there is not 2

OpenStudy (mjdennis):

The problem asks for "point-slope form". All we do is use algebra to move around the pieces of the formula. Here it is with variables instead of numbers:\[m = \frac{ (y - y _{1}) }{ (x - x _{1}) }\] \[m (x - x _{1}) = (y - y _{1})\]

OpenStudy (mellamokatie):

So (8, 3); m = 6 goes into 6(8-8) = (3 - 3)??

OpenStudy (mjdennis):

You said "Wait there is only (8, 3) there is not 2" Correct. The formula has five missing items: m, x, x1, y, y1 We know three things: m, x1, and y1 We are going to leave 'x' and 'y' as 'x' and 'y' for two reasons. A: All the answers have x and y in them, and B: EVERY equation of a LINE needs x and y

OpenStudy (mjdennis):

Does it make sense now?

OpenStudy (mellamokatie):

Yes it does I think thank you!

OpenStudy (mjdennis):

So, what was your answer?

OpenStudy (mellamokatie):

B?

OpenStudy (mellamokatie):

@mjdennis

OpenStudy (mellamokatie):

Hey! I just need someone to say yes B or no B

OpenStudy (mellamokatie):

@rebeccaxhawaii Are you there?

OpenStudy (mjdennis):

You got it.

OpenStudy (mjdennis):

You can tell if I am online if you hover over my name on any post